OpenStudy (anonymous):

If 20a+20b=70, what is the average(arithmetic mean) of a and b?

OpenStudy (anonymous):

Another DBZ fan!!!!!!

hartnn (hartnn):

arithmetic mean will be (a+b)/2 from your equation a+b=7/2 so mean = 7/4 and yes :)

OpenStudy (anonymous):

so wait

OpenStudy (anonymous):

a+b=7/2

OpenStudy (anonymous):

a+b=3.5

hartnn (hartnn):

thats right .

OpenStudy (anonymous):

so the average is 1.75 or 3.5?

hartnn (hartnn):

1.75 because mean = sum/2

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

ty, Son Goku

hartnn (hartnn):

welcome :)